# Triple DES AES Advanced Encryption Standard

### v2

@@ 7.1.1. Existing Attacks against Intel SGX
Given that the essential GPU resource management functions within the GPU driver reside within the TEE enclave, the majority of attacks outlined in Section 6.1 could potentially target this enclave to glean secrets (cf. attack surface ➀ in Figure 5). It is worth noting that while some research may assert that most side-channel attacks against SGX are out of scope, it is critical to provide the corresponding protection when considering the deployment of GPU TEE designs in real-world cloud platforms (where even some forms of physical access by malicious insiders might be relevant). These potential leakages can be categorized as follows: (i) Observing memory access patterns allows attackers to deduce the control flow of programs within the SGX enclave, including decision-making processes. Armed with this insight, attackers could monitor GPU commands and data transfers, thereby inferring the tasks executed on the GPU. For instance, it might become feasible to infer the quantity (and size) of detected objects in each frame of a CNN-based video analysis task (Poddar et al., 2020). (ii) Monitoring page fault patterns enables attackers to extract cryptographic keys from implementations of certain cryptographic open-source software libraries (e.g., OpenSSL and Libgcrypt) (Shinde et al., 2016). With this knowledge, attackers could subsequently access all sensitive information transmitted to the GPU. (iii) Exploiting timing channels allows attackers to infer sensitive information about kernels running on the GPU. For instance, Telekine (Hunt et al., 2020) demonstrates how attackers can accurately classify images from ImageNet by solely observing the timing of GPU kernel execution, without needing access to the images themselves.
@@ 7.1.2. DVFS-based Attacks
On the GPU side. Due to the absence of protection on GPU hardware, attackers with physical access to GPU hardware could directly manipulate GPU voltage, injecting faults into running GPU kernels. For instance, “VoltPillager” (Chen et al., 2021) can inject messages on the SVID bus between the CPU and the voltage regulator on the motherboard, enabling precise control over CPU core voltage. Ultimately, this jeopardizes the confidentiality and integrity of Intel SGX enclaves. Exploring whether attackers can inject malicious SVID packets (or similar) into the respective bus on the GPU card presents an intriguing avenue for further investigation.
@@ 7.1.3. Physical Attacks on PCIe Bus and GPU Memory
In many existing GPU TEEs prototypes lacking vendor hardware support (cf. attack surfaces ➂ and ➄ in Figure 5), GPUs lack a trusted memory region, leaving data in GPU memory exposed in plaintext. Consequently, direct physical access to GPU memory can compromise sensitive data. Further, cold boot attacks (Halderman et al., [n. d.]) could be relevant, if an adversary can ensure (e.g., by cooling) that TEE-protected memory contents are retained and can be readout after a reset of the GPU.
@@ 7.2.1. Existing Attacks against VM-based TEEs
Perhaps because VM-based TEEs have only been proposed in recent years, there is currently limited research on attacks against them. However, recent developments have shown that software-based fault injection attacks can be highly effective against AMD SEV-SNP. These attacks would allow an adversary to compromise the GPU driver within the VM, enabling precise control over sensitive tasks running on the GPU (cf. attack surface ➀ in Figure 6). For example, the “WeSee” attack (Schlüter et al., 2024a) targets AMD SEV-SNP used in NVIDIA H100-based GPU TEEs, allowing an attacker to perform arbitrary memory writes by faking a #$\mathtt{\#VC}$ for MMIO reads. This capability would enabls the attacker to modify commands and data sent to the GPU. More critically, attacks like CacheWarp (Zhang et al., 2024), which remain unmitigated in the latest SEV-SNP implementation, do not depend on the specifics of the guest VM. Therefore, it is crucial to equip NVIDIA H100-based GPU TEEs with protection against these attacks to preserve the confidentiality and integrity of GPU data. It is also important to explore potential attacks against other VM-based TEEs, such as Intel TDX, as these are integral to NVIDIA H100-based GPU TEEs.
@@ 7.2.2. DVFS-based Attacks
On the GPU side. The H100 whitepaper (NVIDIA, 2022b) lacks detail about voltage and frequency management. Assuming an attacker has physical access, they could potentially execute attacks like “VoltPillager” (Chen et al., 2021) to inject malicious SVID packets or equivalent into the relevant GPU card’s bus, thereby altering the GPU voltage.
@@ 7.2.3. Microarchitectural Side-Channel and Covert-Channel Attacks on GPU
The H100 whitepaper (NVIDIA, 2022b) emphasizes the enhanced security features of the Hopper architecture for confidential computing, particularly in collaborative multi-party computing scenarios, namely single tenant with multiple GPUs and multiple tenants with a single GPU. However, vulnerabilities discussed in Section 5.2 may potentially lead to security breaches on the NVIDIA H100 GPU.
[…]
@@ 7.3. Potential Attacks against Arm-based GPU TEE Prototype
Due to various architectural disparities in TEEs for Arm GPUs, several studies have explored Arm-based GPU TEEs. In Figure 7, we base it on the design proposed by StrongBox (Deng et al., 2022). StrongBox serves as a suitable foundation as it was recently introduced and does not rely on specific Arm endpoint features or necessitate hardware modifications to GPU or CPU chips. Notably, while secure GPU computing on Arm or Intel platforms may not align with Arm’s CCA realm-style architecture (Wang et al., 2024b), we omit insights regarding this, as it closely resembles that of NVIDIA Hopper H100-based GPU TEE. In Figure 7, the attacker gains control over the kernel and the entire GPU software stack, including the GPU driver and runtime. This control enables them to manipulate sensitive data and code within GPU applications. They can directly access unified memory used for GPU tasks or manipulate peripherals to evade detection. Furthermore, by compromising the GPU driver, the attacker can undermine the memory management of GPU applications, potentially mapping sensitive data to unprotected regions. Similar to existing x86-based GPU TEEs, trust is placed in the GPU and the secure world. Based on this threat model, we discuss potential attacks in the following.
@@ 7.3.1. Existing Attacks against Arm TrustZone
The attacks summarized in Section 6.2 that target TrustZone are also applicable to this Arm-based GPU TEE prototype. The impact of these side-channel attacks on TrustZone indirectly affects Arm-based GPU TEE (cf. attack surfaces ➀ and ➁ in Figure 7). The potential consequences can be summarized as follows: (i) An attacker could escalate privileges to root to submit numerous malicious tasks (Beniamini, 2016; Shen, 2015). However, establishing covert channels between malicious and victim tasks is unlikely, as GPU tasks are executed sequentially on Arm endpoint GPUs (Deng et al., 2022). (ii) An attacker may monitor cache activity to extract AES keys from TrustZone (Guanciale et al., 2016; Zhang et al., 2016b; Lipp et al., 2016), thus compromising the confidentiality of data transferred between the CPU and GPU.
@@ 7.3.2. Microarchitectural Side-Channel and Covert-Channel Attacks on Integrated GPU
(ii) The existing threat model of Arm-based GPU TEEs may not effectively mitigate attacks exploiting data-dependent DRAM traffic and cache utilization induced by compression to reveal user pixels in the normal world (Wang et al., 2024a). Further testing is needed to assess the success rate of such attacks on Arm-based GPU TEEs.
@@ 7.3.3. Physical Attacks on GPU Memory
Most existing Arm-based GPU TEE prototypes primarily focus on delivering hardware-enforced isolation to protect sensitive code and data, rather than employing memory encryption mechanisms like Intel SGX’s MEE. Consequently, direct physical access to GPU memory (again also possibly via cold boot attacks (Halderman et al., [n. d.])) could potentially compromise sensitive data (cf. attack surface ➃ in Figure 7). For instance, in StrongBox (Deng et al., 2022), a region in the unified memory defined as Secure Task RAM serves as a fixed and non-secure memory region reserved for confidential GPU applications. Its main purpose is to dynamically allocate memory and establish GPU page table mappings and thus prevent unauthorized access control to this region.

### v3

@@ 6.1.1. Existing Attacks against Intel SGX
Given that the essential GPU resource management functions within the GPU driver reside within the TEE enclave, the majority of attacks outlined in the supplemental material could potentially target this enclave to glean secrets (cf. attack surface ➀ in Figure 6). It is worth noting that while some research may assert that most side-channel attacks against SGX are out of scope, it is critical to provide the corresponding protection when considering the deployment of GPU TEE designs in real-world cloud platforms (where even some forms of physical access by malicious insiders might be relevant). These potential leakages can be categorized as follows: (i) Observing memory access patterns allows attackers to deduce the control flow of programs within the SGX enclave, including decision-making processes. Armed with this insight, attackers could monitor GPU commands and data transfers, thereby inferring the tasks executed on the GPU. For instance, it might become feasible to infer the quantity (and size) of detected objects in each frame of a CNN-based video analysis task (Poddar et al., 2020). (ii) Monitoring page fault patterns enables attackers to extract cryptographic keys from implementations of certain cryptographic open-source software libraries (e.g., OpenSSL and Libgcrypt) (Shinde et al., 2016). With this knowledge, attackers could subsequently access all sensitive information transmitted to the GPU. (iii) Exploiting timing channels allows attackers to infer sensitive information about kernels running on the GPU. For instance, Telekine (Hunt et al., 2020) demonstrates how attackers can accurately classify images from ImageNet by solely observing the timing of GPU kernel execution, without needing access to the images themselves.
@@ 6.1.2. DVFS-based Attacks
On the GPU side. Due to the absence of protection on GPU hardware, attackers with physical access to GPU hardware could directly manipulate GPU voltage, injecting faults into running GPU kernels. For instance, “VoltPillager” (Chen et al., 2021) can inject messages on the Serial Voltage Identification (SVID) bus between the CPU and the voltage regulator on the motherboard, enabling precise control over CPU core voltage. Ultimately, this jeopardizes the confidentiality and integrity of Intel SGX enclaves. Exploring whether attackers can inject malicious SVID packets (or similar) into the respective bus on the GPU card presents an intriguing avenue for further investigation.
@@ 6.1.3. Physical Attacks on PCIe Bus and GPU Memory
In many existing GPU TEEs prototypes lacking vendor hardware support (cf. attack surfaces ➂ and ➄ in Figure 6), GPUs lack a trusted memory region, leaving data in GPU memory exposed in plaintext. Consequently, direct physical access to GPU memory can compromise sensitive data. Further, cold boot attacks (Halderman et al., ) could be relevant, if an adversary can ensure (e.g., by cooling) that TEE-protected memory contents are retained and can be readout after a reset of the GPU.
@@ 6.2.1. Existing Attacks against VM-based TEEs
Perhaps because VM-based TEEs have only been proposed in recent years, there is currently limited research on attacks against them. However, recent developments have shown that software-based fault injection attacks can be highly effective against AMD SEV-SNP. These attacks would allow an adversary to compromise the GPU driver within the VM, enabling precise control over sensitive tasks running on the GPU (cf. attack surface ➀ in Figure 7). For example, the “WeSee” attack (Schlüter et al., 2024a) targets AMD SEV-SNP used in NVIDIA H100-based GPU TEEs, allowing an attacker to perform arbitrary memory writes by faking a $\mathtt{\#VC}$ for MMIO reads. This capability would enabls the attacker to modify commands and data sent to the GPU. More critically, attacks like CacheWarp (Zhang et al., 2024), which remain unmitigated in the latest SEV-SNP implementation, do not depend on the specifics of the guest VM. Therefore, it is crucial to equip NVIDIA H100-based GPU TEEs with protection against these attacks to preserve the confidentiality and integrity of GPU data. It is also important to explore potential attacks against other VM-based TEEs, such as Intel TDX, as these are integral to NVIDIA H100-based GPU TEEs.
@@ 6.2.2. DVFS-based Attacks
On the GPU side. The H100 whitepaper (NVIDIA, 2022c) lacks detail about voltage and frequency management. Assuming an attacker has physical access, they could potentially execute attacks like “VoltPillager” (Chen et al., 2021) to inject malicious SVID packets or equivalent into the relevant GPU card’s bus, thereby altering the GPU voltage.
@@ 6.2.3. Microarchitectural Side-Channel and Covert-Channel Attacks on GPU
The H100 whitepaper (NVIDIA, 2022c) emphasizes the enhanced security features of the Hopper architecture for confidential computing, particularly in collaborative multi-party computing scenarios, namely single tenant with multiple GPUs and multiple tenants with a single GPU. However, vulnerabilities discussed in Section 5.2 may potentially lead to security breaches on the NVIDIA H100 GPU.
Single tenant with multiple GPUs. In the scenario of single tenant with multiple GPUs, a single VM is concurrently assigned multiple physically independent GPUs, each connected via NVLink. Recall that Dutta et al. (Dutta et al., 2023) demonstrate that an attacker on one GPU could induce contention on the L2 cache of another GPU via NVLink, enabling a covert channel and Prime+Probe attacks across GPUs. However, such attacks may face challenges in the NVIDIA H100 GPU model due to two key factors (cf. attack surface ➁ in Figure 7). Firstly, the attacker’s process needs access to a remote GPU’s memory through NVLink, which may be restricted in the threat model of the NVIDIA H100 GPU TEE. Thus, the challenge lies in how to enable the spy process to run within one of the assigned GPUs and have access to another GPU’s memory. Secondly, the H100 GPU features a 50 MB L2 cache, $1.25{\times}$ larger than the A100 GPU. Nevertheless, a larger L2 cache with more sets and ways can make Prime+Probe attacks more difficult, because the attacker needs to fill and monitor a larger number of cache lines.
@@ 6.3. Potential Attacks against Arm-based GPU TEE Prototype
Due to various architectural disparities in TEEs for Arm GPUs, several studies have explored Arm-based GPU TEEs. In Figure 8, we base it on the design proposed by StrongBox (Deng et al., 2022). StrongBox serves as a suitable foundation as it was recently introduced and does not rely on specific Arm endpoint features or necessitate hardware modifications to GPU or CPU chips. Notably, while secure GPU computing on Arm or Intel platforms may not align with Arm’s CCA realm-style architecture (Wang et al., 2024a), we omit insights regarding this, as it closely resembles that of NVIDIA Hopper H100-based GPU TEE. In Figure 8, the attacker gains control over the kernel and the entire GPU software stack, including the GPU driver and runtime. This control enables them to manipulate sensitive data and code within GPU applications. They can directly access unified memory used for GPU tasks or manipulate peripherals to evade detection. Furthermore, by compromising the GPU driver, the attacker can undermine the memory management of GPU applications, potentially mapping sensitive data to unprotected regions. Similar to existing x86-based GPU TEEs, trust is placed in the GPU and the secure world. Based on this threat model, we discuss potential attacks in the following.
@@ 6.3.1. Existing Attacks against Arm TrustZone
The attacks summarized in the supplemental material that target TrustZone are also applicable to this Arm-based GPU TEE prototype. The impact of these side-channel attacks on TrustZone indirectly affects Arm-based GPU TEE (cf. attack surfaces ➀ and ➁ in Figure 8). The potential consequences can be summarized as follows: (i) An attacker could escalate privileges to root to submit numerous malicious tasks (Beniamini, 2016; Shen, 2015). However, establishing covert channels between malicious and victim tasks is unlikely, as GPU tasks are executed sequentially on Arm endpoint GPUs (Deng et al., 2022). (ii) An attacker may monitor cache activity to extract AES keys from TrustZone (Guanciale et al., 2016; Zhang et al., 2016c; Lipp et al., 2016), thus compromising the confidentiality of data transferred between the CPU and GPU.
@@ 6.3.2. Microarchitectural Side-Channel and Covert-Channel Attacks on Integrated GPU
(ii) The existing threat model of Arm-based GPU TEEs may not effectively mitigate attacks exploiting data-dependent DRAM traffic and cache utilization induced by compression to reveal user pixels in the normal world (Wang et al., 2024b). Further testing is needed to assess the success rate of such attacks on Arm-based GPU TEEs.
@@ 6.3.3. Physical Attacks on GPU Memory
Most existing Arm-based GPU TEE prototypes primarily focus on delivering hardware-enforced isolation to protect sensitive code and data, rather than employing memory encryption mechanisms like Intel SGX’s MEE. Consequently, direct physical access to GPU memory (again also possibly via cold boot attacks (Halderman et al., )) could potentially compromise sensitive data (cf. attack surface ➃ in Figure 8). For instance, in StrongBox (Deng et al., 2022), a region in the unified memory defined as Secure Task RAM serves as a fixed and non-secure memory region reserved for confidential GPU applications. Its main purpose is to dynamically allocate memory and establish GPU page table mappings and thus prevent unauthorized access control to this region.
@@ 6.3.4. DVFS-based Attacks
6.4. Comparison of Attacks on Different GPU TEEs
x86-based GPU TEE prototypes, NVIDIA Hopper H100, and Arm-based GPU TEEs adopt distinct trust models, hardware assumptions, and protection strategies. x86-based designs (e.g., Graviton, HIX, HETEE) typically build on Intel SGX, placing only essential GPU driver components inside the enclave to minimize the TCB. These prototypes often lack GPU-side memory encryption and leave out DVFS and power management from protection, making them susceptible to software-based fault injection and physical attacks (e.g., cold boot, SVID manipulation). In contrast, NVIDIA Hopper H100 integrates vendor-supported confidential computing features, including memory encryption and compatibility with VM-based TEEs. While this offers system-wide isolation, it inherits known vulnerabilities of VM-based TEEs, such as software-based fault injection (e.g., WeSee) and cache manipulation (e.g., CacheWarp). Furthermore, the limited public documentation presents challenges for third-party validation and systematic defense design. Arm-based GPU TEE designs, such as StrongBox, use TrustZone to provide software-based isolation without requiring hardware modifications. This approach improves portability and ease of deployment but sacrifices granularity. These systems typically lack hardware-enforced memory encryption and rely on unified memory and shared caches, which can expose sensitive data to physical and side-channel attacks. Operating in resource-constrained environments (e.g., edge devices), Arm-based TEEs tend to have a smaller TCB but may not adequately defend against DVFS abuse or microarchitectural leakage.
In summary, existing x86-based prototypes prioritize minimal hardware changes but depend on SGX’s integrity; NVIDIA H100 adopts holistic, hardware-enhanced isolation but inherits the complexity and weaknesses of VM-based TEEs; and Arm-based TEEs favor lightweight integration at the cost of fine-grained protection. From an attack perspective, x86 prototypes are vulnerable at the interface between CPU TEEs and unprotected GPUs, NVIDIA’s model is exposed through VM-level fault channels and microarchitectural sharing, and Arm designs are weakest at the hardware and cache/memory system level. These architectural differences reflect distinct trade-offs and highlight the need for architecture-aware defenses tailored to each GPU TEE design.
@@ Acknowledgements
Appendix A Security of CPU TEEs
In this online appendix, we introduce typical attacks against CPU TEEs. As there already exist survey papers focusing on reviewing existing attacks on CPU TEEs, and the aim of this survey is to explore potential attacks on GPU TEEs, we simply categorize the most common attacks based on their technical details and existing surveys (Fei et al., 2021; Nilsson et al., 2020; Cerdeira et al., 2020; Li, 2022) here. We provide a brief summary of these attacks regarding attack surfaces and countermeasures in table 4.
Table 4. Summary of attacks on CPU TEEs
CPU TEE Attacks Ref Attack Surfaces Countermeasures
Arch. Threat Model PT SE $C$ DRAM BTB PHT RSB IFC HC Bug INT
Trusted Untrusted
SGX CPU package, EPC, MEE OS/Kernel AT (Xu et al., 2015) Obfuscation, randomization (Brasser et al., 2019a; Ahmad and Bates, 2019)
(Van Bulck et al., 2017b; Kim et al., 2019)
(Gyselinck et al., 2018)
Cache (Brasser et al., 2017b; Dall et al., 2018; Götzfried et al., 2017; Moghimi et al., 2017) Obfuscation, randomization, anomaly detection (Brasser et al., 2019a; Ahmad and Bates, 2019)
(Schwarz et al., 2017)
BP (Lee et al., 2017) Obfuscation (Ahmad and Bates, 2019)
(Chen et al., 2019)
(Evtyushkin et al., 2018; Huo et al., 2020)
(Koruyeh et al., 2018)
(Van Bulck et al., 2018)
TE (Van Bulck et al., 2018) Microcode updates from Intel
(Van Schaik et al., 2020; Ragab et al., 2021)
SwFIA (Murdock et al., 2020; Kenjar et al., 2020; Qiu et al., 2020) Microcode updates from Intel, restricted access to registers, ECC
(Lipp et al., 2021; Wang et al., 2022b; Liu et al., 2022)
(Jang et al., 2017)
Hw (Chen et al., 2021; Chen and Oswald, 2023; Lee et al., 2020) Fuse-level lockouts, access pattern obfuscation, harden BMC/PMU firmware
Interface (Van Bulck et al., 2019; Alder et al., 2020; Cloosters et al., 2020; Khandaker et al., 2020; Alder et al., 2024) State and register sanitization, minimize the ABI surface, automated runtime analysis
TrustZone SW, Secure monitor NW OS IB (Beniamini, 2016; Shen, 2015; Komaromy, 2018) Software-only verification (Ferraiuolo et al., 2017; Jin et al., 2022)
Cache (Zhang et al., 2016a; Guanciale et al., 2016; Lipp et al., 2016; Zhang et al., 2016c; Cho et al., 2018) Cache isolation (Dessouky et al., 2021)
BP (Ryan, 2019) Flush BTB
SwFIA (Tang et al., 2017; Qiu et al., 2020; Carru, 2017) Restrict access to DVFS interfaces
SEV PSP, MEE OS/ Hypervisor ME (Hetzelt and Buhren, 2017; Buhren et al., 2017; Du et al., 2017; Morbitzer et al., 2018; Wilke et al., 2020; Li et al., 2021b) Mitigated in AMD SEV-SNP
(Morbitzer et al., 2018)
(Li et al., 2021a)
SwFIA (Morbitzer et al., 2021; Zhang et al., 2024) Microcode update from AMD
ID (Schlüter et al., 2024a, b; Wilke et al., 2024) Isolate interrupt handlers, detect single-stepping
Attacks contains Address Translation (AT), Cache, Branch Prediction (BP), Transient Execution (TE), Software-based Fault Injection (SwFIA), Interface, Hardware (Hw), Implementation Bug (IB), Memory Encryption (ME), Interrupt-driven (ID). Attack Surfaces includes Page Table (PT), Segmentation (SE), CPU Cache (CC), Branch Target Buffer (BTB), Page History Table (PHT), Return Stack Buffer (RSB), Interface (IFC, containing DVFS interfaces, enclave interfaces, etc.), Interrupt (INT, interrupt controller/interrupt handling logic).
A.1. Security of Intel SGX
A.1.1. Controlled Channel and Address Translation-based Attacks
Controlled channels attack, a concept introduced by Xu et al. (Xu et al., 2015) in 2015, exploits the privileges of the OS to monitor and manipulate page tables. This allows the attacker to uncover the memory access patterns of the enclave code, cf. also (Shinde et al., 2016). This attack is especially effective because, while Intel SGX enclaves protect sensitive computations from a compromised OS, the OS still manages memory and can observe page faults, thereby inferring details about the enclave’s operations. SGX-Step (Van Bulck et al., 2017a) serves as a practical tool for, among others, executing such controlled channel attacks by configuring Advanced Programmable Interrupt Controller (APIC) timers, issuing interrupts, and tracking page-table entries. SGX-Step enables deterministic single-stepping of enclave execution, allowing for precise monitoring of memory access patterns over time.
A.1.2. Cache-based Attacks
A.1.3. Branch Prediction Attacks
A.1.4. Transient Execution Attacks
Transient execution attacks, which exploit out-of-order and speculative execution of instructions, have also been studied in the context of SGX. The aforementioned BTB and RSB attacks (Chen et al., 2019; Koruyeh et al., 2018) can also fall into this class of attacks, stemming from Spectre (Kocher et al., 2020). Additionally, the Foreshadow attack (Van Bulck et al., 2018), which targets SGX by employing Meltdown-like techniques, breaches enclave execution confidentiality. While MDS attacks (Canella et al., 2019; Van Schaik et al., 2019; Schwarz et al., 2019) also leverage speculative and out-of-order execution, they primarily exploit information leakage from various implementation-specific and undocumented intermediary buffers of the targeted microarchitecture. SGAxe (Van Schaik et al., 2020) and Crosstalk (Ragab et al., 2021) extend MDS to extract private keys from SGX enclaves.
A.1.5. Software-based Fault Injection and Side-Channel Attacks
A.1.6. Hardware-based Attacks
Apart from software-based fault injection attacks, Chen et al. (Chen et al., 2021; Chen and Oswald, 2023) explore hardware-level fault injection attacks on SGX enclaves. “VoltPillager” is a low-cost tool that injects messages on the SVID bus between the CPU and the voltage regulator, breaching the confidentiality and integrity of Intel SGX enclaves (Chen et al., 2021). Chen and Oswald later demonstrates that undervolting via the Power Management Bus (PMBus) can similarly break SGX enclave integrity, bypassing Intel’s countermeasures against previous software undervolting attacks (Murdock et al., 2020; Kenjar et al., 2020; Qiu et al., 2020). “MEMBUSTER” (Lee et al., 2020) introduces an off-chip attack that breaks SGX enclave confidentiality by snooping the memory bus (with high-spec, expensive lab equipment), extracting memory access patterns and increasing cache misses using techniques like critical page whitelisting and cache squeezing.
A.1.7. Interface-based Attacks
A.1.8. Countermeasures
(iii) Randomization: “SGX-Shield” (Seo et al., 2017) leverages ASLR to create a secure in-enclave loader that secretly randomizes memory space layout. Hosseinzadeh et al. (Hosseinzadeh et al., 2018) propose runtime control flow randomization, implemented as compiler extensions on llvm. “DR.SGX” (Brasser et al., 2019a) disrupts memory observations by permuting data locations.
(iv) Application/Source code design: Early works like “OBFUSCURO” (Ahmad and Bates, 2019) enforce operations with Oblivious RAM (ORAM) protection to prevent attackers from inferring access patterns. Several schemes (Ohrimenko et al., 2016; Poddar et al., 2020; Wang et al., 2022a) utilize oblivious primitives, implemented in x86_64 assembly, to design data-oblivious applications.
A.2. Security of Arm TrustZone
A.2.1. Implementation Bugs
Cerdeira et al. (Cerdeira et al., 2020) examine bug reports obtained from public Common Vulnerabilities and Exposures (CVE) databases and vendor bulletin reports, categorising them into validation bugs, functional bugs, and extrinsic bugs. These bugs are pervasive and often exploited to escalate privileges, enabling attackers to fully hijack kernels in devices from Qualcomm (Beniamini, 2016) and Huawei (Shen, 2015) or compromise client applications like Samsung Pay (Komaromy, 2018).
A.2.2. Cache-based Attacks
Given that the cache is shared between NW and SW in TrustZone-enabled processors, several studies capitalize on the contention arising from this cache coherency to extract sensitive information (Zhang et al., 2016a; Guanciale et al., 2016; Lipp et al., 2016; Zhang et al., 2016c; Cho et al., 2018). (Lipp et al., 2016; Zhang et al., 2016c; Cho et al., 2018) use the Prime+Probe method to monitor the cache activities within TrustZone.
A.2.3. Branch Prediction Attacks
A.2.4. Software-based Fault Injection Attacks
A.2.5. Countermeasures
In this section, we review and categorize countermeasures based on both previous survey (Cerdeira et al., 2020) and potential solutions from the summarized attack papers. In addressing the architectural defences, some research efforts (Sun et al., 2015; Brasser et al., 2019b) enhance isolation granularity between TEE components and reduce the amount of code running in SW, thereby limiting severe privilege escalation attacks within SW. Countermeasures (Jang et al., 2015; Jang and Kang, 2018) tackle the weaknesses of inadequate/weak authentication when accessing TEE resources from NW, and potentially insecure shared memory used for data exchange across the NW-SW boundary. Unlike Intel SGX, TrustZone lacks built-in on-chip memory encryption, prompting some works (Zhang et al., 2016b; Yun and Zhong, 2019) to design encryption strategies to bridge this gap. To further strengthen the verification of both TEE and TA binaries’ integrity and identity, several studies (Ferraiuolo et al., 2017; Lentz et al., 2018) introduced additional primitives, such as enhanced remote attestation and sealed storage. To address implementation bugs, “RustZone” (Evenchick, 2018) extends TrustZone by implementing TAs using Rust, which inherently provides memory safety and thread safety.
Cache side-channel attacks can be mitigated by carefully implementing cryptographic algorithms in software (Guanciale et al., 2016) or using dedicated hardware (Lipp et al., 2016). In terms of BTB attacks, flushing the shared microarchitectural structures when transferring between NW and SW can prevent attacks from targeting keys through the BTB. However, as observed in the Nexus 5X firmware by (Ryan, 2019), the L1 cache is not flushed, making Prime+Probe-based attacks still feasible.
A.3. Security of AMD SEV
A.3.1. Memory Encryption Issues
A.3.2. Software-based Fault Injection Attacks
Morbitzer et al. (Morbitzer et al., 2021) exploit the system’s lack of memory integrity protection to inject faults and execute arbitrary code within SEV VMs through I/O channels, using the hypervisor to locate and trigger the execution of encrypted payloads. However, SEV-SNP (Sev-Snp, 2020) can mitigate this issue through Reverse Map Table (RMP). Buhren et al. (Buhren et al., 2021) present a voltage fault injection attack enabling the execution of custom payloads on AMD secure processors, subsequently decrypting a VM’s memory by deploying custom SEV firmware. “CacheWarp” (Zhang et al., 2024) presents a new software-based fault attack targeting AMD SEV-ES and SEV-SNP, exploiting the possibility to revert modified cache lines of guest VMs to their previous (stale) state at the architectural level. This attack only requires interrupting the VM at a point chosen by the attacker to invalidate modified cache lines without writing them back to memory.
A.3.3. Interrupt-driven Attacks
Similar to SGX-Step framework (Van Bulck et al., 2017a), SEV-Step (Wilke et al., 2024) introduces reliable single-stepping against SEV VMs. In addition to this functionality, SEV-Step facilitates access to common attack techniques, such as page fault tracking and cache attacks targeting SEV. More recently, Schlüter et al. (Schlüter et al., 2024a) introduce a novel attack where the hypervisor injects well-crafted, malicious $\mathtt{\#VC}$ (a new exception by AMD facilitating communication between the VM and untrusted hypervisor) into a victim VM’s CPU, enabling the attacker to induce arbitrary behaviour in the VM. In another related attack, “HECKLER” (Schlüter et al., 2024b) shows that the hypervisor can inject malicious non-timer interrupts to compromise VMs. They leverage interrupt handlers with global effects, allowing manipulation of a VM’s register states to alter data and control flow. In general, SEV-SNP remains vulnerable to these attacks.
A.3.4. Countermeasures